# SFNet: A Spatio-Frequency Domain Deep Learning Network for Efficient Alzheimer’s Disease Diagnosis

### v1

@@ Abstract
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that predominantly affects the elderly population and currently has no cure. Magnetic Resonance Imaging (MRI), as a non-invasive imaging technique, is essential for the early diagnosis of AD. MRI inherently contains both spatial and frequency information, as raw signals are acquired in the frequency domain and reconstructed into spatial images via the Fourier transform. However, most existing AD diagnostic models extract features from a single domain, limiting their capacity to fully capture the complex neuroimaging characteristics of the disease. While some studies have combined spatial and frequency information, they are mostly confined to 2D MRI, leaving the potential of dual-domain analysis in 3D MRI unexplored. To overcome this limitation, we propose Spatio-Frequency Network (SFNet), the first end-to-end deep learning framework that simultaneously leverages spatial and frequency domain information to enhance 3D MRI-based AD diagnosis. SFNet integrates an enhanced dense convolutional network to extract local spatial features and a global frequency module to capture global frequency-domain representations. Additionally, a novel multi-scale attention module is proposed to further refine spatial feature extraction. Experiments on the Alzheimer’s Disease Neuroimaging Initiative (ANDI) dataset demonstrate that SFNet outperforms existing baselines and reduces computational overhead in classifying cognitively normal (CN) and AD, achieving an accuracy of 95.1%.
@@
Alzheimer’s Disease, Spatio-Frequency Domain, Attention Mechanism, Multi-Scale Features, Deep Learning
@@ I INTRODUCTION
Despite significant progress achieved by frequency domain models, several limitations remain. First, certain models, such as GFNet, exclusively extracted features from the frequency domain while neglecting the spatial structural information inherent in MRI data. Second, although some models integrated spatial and frequency domain features, they often rely on 2D slices, resulting in a lack of spatial continuity across the three-dimensional volume. To overcome these limitations, we propose a 3D deep learning model that integrates spatial and frequency domain features to fully exploit the complementary information embedded in MRI data. In the spatial domain, an improved DenseNet architecture is employed to extract local structural features, enhanced by a multi-scale spatial attention module to improve the perception of brain regions at different spatial resolutions and scales. In the frequency domain, a global frequency module applies the fast Fourier transform to feature maps, capturing long-range dependencies across regions and enabling effective global context modeling. By fusing spatio-frequency domain features, the model enhances inter-regional interactions and significantly improves its discriminative performance in predicting early-stage Alzheimer’s disease. The main contributions of this paper are as follows:
  • We propose a novel AD diagnosis model named (Spatio-Frequency Network) SFNet that the first to integrate spatial local features and global dependencies of frequency domain based on 3D MRI, enhancing the accuracy of AD classification.
@@ II-B Image Preprocessing
Figure 2: SFNet architecture
@@ III-A Spatial Feature Extraction Module
The spatial feature extraction modul consists of two components. First, DenseNet is employed as the backbone network to extract spatial features. Then, a novel multi-scale attention module is incorporated to progressively expand the receptive field and capture multi-scale local patterns.
@@ IV-B Model Comparsion On All Three Classification Tasks
In this section, we compare the proposed SFNet with recent related methods for the classification of AD/CN, CN/MCI, and AD/MCI, as summarized in Table II. The results show that SFNet outperforms GFNet, highlighting the effectiveness of incorporating spatio-frequency domain information for feature extraction. Compared to Transformer-based models, such as ViT, as well as CNN-Transformer-based models like M3T and CViT, SFNet consistently demonstrates superior performance. Specifically, it achieves higher ACC, SEN, and AUC while maintaining a lower number of parameters. This superiority stems from SFNet’s ability to capture long-range dependencies in the frequency domain, which more effectively complements spatial features than the spatial self-attention mechanisms used in CNN-Transformer-based models. Notably, compared to the best-performing Conv-Swinformer method, SFNet reduces model complexity and achieves improvements of 1.6% in AD/CN classification accuracy, 6.9% in CN/MCI, and 2.8% in AD/MCI.
@@ V Conclusions
In this paper, we propose an end-to-end deep learning framework, SFNet, for the early diagnosis of Alzheimer’s disease using 3D MRI data. The model incorporates a multiscale attention module and DenseNet for spatial feature extraction, alongside the global frequency module to capture frequency-domain information. This dual-domain approach enables SFNet to effectively learn both local fine-grained textures and global structural patterns. Experiments on ANDI datasets demonstrate that SFNet outperforms recent state-of-the-art methods in classification accuracy while maintaining a lower total parameters. Furthermore, we visualize the learnable global filters in the frequency domain that reveal how SFNet progressively adapts its attention to different spectral components, contributing to the interpretability of the model. In future work, we plan to focus on multi-model integration to enhance the robustness and generalizability of SFNet.

### v2

# SFNet: A Spatial-Frequency Domain Deep Learning Network for Efficient Alzheimer’s Disease Diagnosis

@@ Abstract
Alzheimer’s disease (AD) is a progressive neurodegenerative disorder that predominantly affects the elderly population and currently has no cure. Magnetic Resonance Imaging (MRI), as a non-invasive imaging technique, is essential for the early diagnosis of AD. MRI inherently contains both spatial and frequency information, as raw signals are acquired in the frequency domain and reconstructed into spatial images via the Fourier transform. However, most existing AD diagnostic models extract features from a single domain, limiting their capacity to fully capture the complex neuroimaging characteristics of the disease. While some studies have combined spatial and frequency information, they are mostly confined to 2D MRI, leaving the potential of dual-domain analysis in 3D MRI unexplored. To overcome this limitation, we propose Spatial-Frequency Network (SFNet), the first end-to-end deep learning framework that simultaneously leverages spatial and frequency domain information to enhance 3D MRI-based AD diagnosis. SFNet integrates an enhanced dense convolutional network to extract local spatial features and a global frequency module to capture global frequency-domain representations. Additionally, a novel multi-scale attention module is proposed to further refine spatial feature extraction. Experiments on the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset demonstrate that SFNet outperforms existing baselines and reduces computational overhead in classifying cognitively normal (CN) and AD, achieving an accuracy of 95.1%.
@@
Alzheimer’s Disease, Spatial-Frequency Domain, Attention Mechanism, Multi-Scale Features, Deep Learning
@@ I INTRODUCTION
Despite significant progress achieved by frequency domain models, several limitations remain. First, certain models, such as GFNet, exclusively extracted features from the frequency domain while neglecting the spatial structural information inherent in MRI data. Second, although some models integrated spatial and frequency domain features, they often rely on 2D slices, resulting in a lack of spatial continuity across the three-dimensional volume. To overcome these limitations, we propose a 3D deep learning model that integrates spatial and frequency domain features to fully exploit the complementary information embedded in MRI data. In the spatial domain, an improved DenseNet architecture is employed to extract local structural features, enhanced by a multi-scale spatial attention module to improve the perception of brain regions at different spatial resolutions and scales. In the frequency domain, a global frequency module applies the fast Fourier transform to feature maps, capturing long-range dependencies across regions and enabling effective global context modeling. By fusing spatial-frequency domain features, the model enhances inter-regional interactions and significantly improves its discriminative performance in predicting early-stage Alzheimer’s disease. The main contributions of this paper are as follows:
  • We propose a novel model for AD diagnosis, termed SFNet, which is the first to integrate spatial local features and global frequency-domain dependencies based on 3D MRI data, effectively enhancing the accuracy of AD classification.
@@ II-B Image Preprocessing
Figure 2: The architecture of SFNet
@@ III-A Spatial Feature Extraction Module
The spatial feature extraction module consists of two components. First, DenseNet is employed as the backbone network to extract spatial features. Then, a novel multi-scale attention module is incorporated to progressively expand the receptive field and capture multi-scale local patterns.
@@ IV-B Model Comparsion On All Three Classification Tasks
In this section, we compare the proposed SFNet with recent related methods for the classification of AD/CN, CN/MCI, and AD/MCI, as summarized in Table II. The results show that SFNet outperforms GFNet, highlighting the effectiveness of incorporating spatial-frequency domain information for feature extraction. Compared to Transformer-based models, such as ViT, as well as CNN-Transformer-based models like M3T and CViT, SFNet consistently demonstrates superior performance. Specifically, it achieves higher ACC, SEN, and AUC while maintaining a lower number of parameters. This superiority stems from SFNet’s ability to capture long-range dependencies in the frequency domain, which more effectively complements spatial features than the spatial self-attention mechanisms used in CNN-Transformer-based models. Notably, compared to the best-performing Conv-Swinformer method, SFNet reduces model complexity and achieves improvements of 1.6% in AD/CN classification accuracy, 6.9% in CN/MCI, and 2.8% in AD/MCI.
@@ V Conclusions
In this paper, we propose an end-to-end deep learning framework, SFNet, for the early diagnosis of Alzheimer’s disease using 3D MRI data. The model incorporates a multiscale attention module and DenseNet for spatial feature extraction, alongside the global frequency module to capture frequency-domain information. This dual-domain approach enables SFNet to effectively learn both local fine-grained textures and global structural patterns. Experiments on ADNI datasets demonstrate that SFNet outperforms recent state-of-the-art methods in classification accuracy while maintaining a lower total parameters. Furthermore, we visualize the learnable global filters in the frequency domain that reveal how SFNet progressively adapts its attention to different spectral components, contributing to the interpretability of the model. In future work, we plan to focus on multi-model integration to enhance the robustness and generalizability of SFNet.